Muon Spectrometer White Board

Muon ATLAS Control Room 71363 & 71365
Muon Satellite Control Room 70948
Muon Run Coordinators: Stephanie Zimmermann (163050) and Enrico Pasqualucci (162197)

On-call Experts:

ATLAS Run plan

Last runs in the ATLAS partition (under development)



Last page update 04 Dec 2020 06:46:54

Additional Muon Plans and Muon Run Coordinator's notes

Notes:

  • Please check the muon calibration stream to be working/running when stable beams are reached.
  • TGC histograms have been removed for the time being from OHP (performance problems).

Shift Message Board

Shifters and Experts, when posting a comment please use a bullet (on a new line, type three spaces and then a *). Please add a date and your name so that we are aware that the information is recent

The Known Problems part in this section can be updated by everybody.

MDT Known Problems and Issues
  • 2011 - Missing MDT EE chambers: EE chambers other than EEL sector 5 A+C, EEL sector 13 A+C, EEL sector 11 side A, and EEL sector 15 side C are not yet installed and thus excluded or absent from the DCS FSM and disabled from readout in the MDT Control panel.

CSC Known Problems and Issues
  • 2011 - No HV for Layers AL05 Layer 2 and AL09 Layer 4: HV permanently OFF due to broken wires.
  • 2011 - No DCS temperature monitoring for sectors AL15,AS16 side A, sectors disabled in the DCS FSM (Temperature subtree, and they depend on neighboor chambers. AS16 depends on AL01 and AL15 depends on AS14)

RPC Known Problems and Issues

TGC Known Problems and Issues
  • 2011.04.13 - No TGC histograms in OHP for the moment. TGC histogram configuration is under revision, the old setup crashed OHP (X-windows load issue).
  • 2011.04.17 - Stephanie, Masaya - Bad (too low) voltage for ASD pos and neg M3 sector 02 side A, LV board to be replaced during next tech stop. Alarms on voltages < 2.7 V deactivated in DCS (remember to reactivate after board replaced).

Special Shifter Instructions and Points Of Attention

Expert Message Board

Experts, feel free to put items currently to be watched out for or under investigation here. Please include your name and a date.

MDT Issues to watch out for ....

  • 2011.02.16 - Stephanie - Temperature for BrCtrl 14 Crate 0 Boards slightly high, the slow temperature increase also observed halted after tech stop end March and increase in cooling loop pressure. If it's stable, keep it.
  • 2011.04.04 - Stephanie - BIS6A08 Mezz 0x200 disabled, persistent chamber being dropped from the run. BIS1C16 mezz 0x800 disabled, chamber dropped frequently and errors in L1 order. Check next tech stop.
  • 2011.04.05 - - Bad projective alignment line in sector 1, check next tech stop.
  • 2011.04.04 - Stephanie - Intermittent MDT PS 48V Y2505S2 Generator1 FanStatus ABNORMAL Warning in DCS alarm screen, voltage jump from 3.9 to 4.5 V. To be investigated.

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2011-04-17 - unknown
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Sandbox All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback